# Charged superradiant instability in a spherical symmetric regular black hole

### v1

@@ Abstract
We investigate the stability of a spherically symmetric regular black hole under charged scalar field perturbations. The findings reveal the presence of charged superradiant instability within specific parameter ranges, with the growth rate potentially reaching a high order of magnitude up to approximately $\text{Im}(M\omega)\sim 10^{-3}$. This value significantly surpasses the instability observed in ABG black holes reported in earlier studies, implying a potentially relevant timescale for observing this phenomenon in astrophysical scenarios. Furthermore, a detailed analysis is presented on how the model’s three parameters affect the onset and intensity of the instability. Our study contributes further insights into the potential manifestation of this instability in spherical regular black holes and its correlation with the nonlinear effects of the electromagnetic field.
@@ I Introduction
Singularities are commonly considered non-physical within the classical theoretical framework due to their infinite energy density, prompting researchers to seek resolution through the inclusion of quantum effects in matter or gravity. While singularity theorems rely on strict conditions like the null energy condition for matter, allowing for the possibility of singularity-free black holes. To address the issue of non-physical singularities, a significant effort is being made by researchers to develop regular black hole solutions. This endeavor dates back to 1968 when Bardeen introduced a regular black hole model that adhered to the weak energy condition [14], circumventing the singularity theorems. Subsequently, Ayon-Beato and Garcia established a class of regular black holes known as ABG black holes [15, 16], featuring spherical charged configurations within the framework of Einstein-nonlinear electromagnetic fields. Further advancements led them to realize the Bardeen black hole as an exact solution within some Einstein-nonlinear-Maxwell theories [17]. These breakthroughs have inspired the development of various regular black hole models predominantly supported by nonlinear electromagnetic fields [18] (see reviews [19, 20, 21] and references therein), which serve as an avenue to incorporate quantum effects in the electromagnetic field theory [22, 23, 24, 25]. Extensive investigations have been conducted on the dynamic behaviors, thermodynamic properties, and astrophysical implications of these black holes. The distinct observational effects exhibited by regular black holes compared to conventional ones have been a subject of interest, as indicated by recent studies and reviews on the topic [19, 20, 21, 26, 27, 28, 29].
@@
The authors of Refs. [30, 31, 32] recently conducted an extensive examination of the superradiance phenomenon in ABG black holes subjected to charged scalar field perturbations. Their findings suggest that within certain parameter ranges, ABG black holes display superradiant instability, a characteristic distinct from the well-established behavior of Reissner-Nordström (RN) black holes. The study of superradiance and its associated instability in black holes has been a focal point of research in black hole physics for a significant period. Superradiant instability typically manifests in rotating black holes [33, 34, 35, 36, 37]. However, for non-rotating, spherically symmetric RN black holes, superradiance can occur under charged scalar field perturbations without leading to instability [38, 39]. The research outcomes in the Refs. [30, 31, 32] introduce a novel scenario: when accounting for the nonlinear effects of the electromagnetic field, spherically symmetric charged regular black holes may also exhibit superradiant instability. The superradiant instability of black holes holds significant astronomical observational implications. This mechanism results in the accumulation of scalar particles around the black hole, giving rise to observable effects such as alterations in gravitational waveforms and interference with the imaging of the black hole shadow [40, 41, 42, 43, 44, 45, 46]. These observable effects offer valuable insights and avenues for exploring exotic particles beyond the standard model and furthering our comprehension of the physical characteristics of black holes. See more discussions on this phenomenon and its astrophysical implications, please refer to the review [47] and references therein. Thus, the research findings in Refs. [30, 31, 32] not only unveil the superradiant instability of ABG black holes under specific circumstances but also present a potential observational approach for identifying and investigating the properties of regular black holes.
@@
The inquiry arises as to whether the phenomenon of superradiant instability is exclusive to ABG black holes or is a common occurrence among regular black holes. This study delves into this question by exploring a distinct class of regular black holes [48] generated through the ”gravitational-decoupling” (GD) technique, possessing unique characteristics. These black holes’ spacetime transitions between Minkowski and Schwarzschild spacetime are governed by a deformation parameter and can also be supported by a nonlinear electromagnetic field. Through a detailed analysis of the behavior of charged scalar field perturbations around such black holes, it is observed that the presence and strength of superradiant instability are closely linked to the deformation parameter. Remarkably, when the deformation parameter falls within a specific range, the intensity of superradiant instability can escalate significantly, surpassing that of ABG black holes. These findings provide evidence that regular black holes typically display superradiant instability.
@@
This paper is structured as follows: Section II provides a concise overview of the standard black hole under consideration. Section III delves into the equation of motion governing the tested charged scalar field. In Section IV, an examination of the parameter space is conducted to determine the permissible region for initiating the superradiant instability. Section V employs a numerical approach to identify the superradiant unstable modes. The concluding section encompasses a summary and discussions. The study employs natural units with $G=c=\hbar=4\pi\epsilon_{0}=1$ throughout, where $G,c,\hbar$ and $\epsilon_{0}$ are the Newtonian gravitational constant, the photon velocity in vacuum, the induced planck constant and the vacuum permittivity.
@@ II The spacetime
where $M$ is the ADM mass of the black hole and $\alpha$ is a deformation parameter. It is constructed through the deformation of the Schwarzschild spacetime with the method of gravitational decoupling. The metric represents a spacetime interpolating the Minkowski spacetime and the Schwarzschild black hole: In the limit of $\alpha\rightarrow 0$, it reduces to the Minkowski metric; while in the limit of $\alpha\rightarrow\infty$, it reduces to the standard Schwarzschild metric. It should also be noted that the deformation of this spacetime with respect to the Schwarzschild metric takes an exponential decaying form, so that the deformation mainly affects the small-$r$ region but leaves the large-$r$ region little affected. The spacetime has a de-Sitter core and is regular at $r=0$ which can be seen by calculating the typical curvature invariants, for example,
@@
It asymptotically takes a Yukawa-like form rather than a Maxwell-like form, which is very different from that of ABG black holes [15, 16]. So the“nonlinear electromagnetic field” may not be considered as the usual nonlinear extension of Maxwell’s theory simply, rather it is more likely associated with some kind of short-range interaction. The action from which to derive this solution is rather involved and not relevant to our present study, so we will not dispaly it here. For more details, one can refer to Ref. [48].
The event horizon $r=r_{h}$ of the black hole is given by the largest root of $f(r)=0$, which depends on the value of $\alpha$. In Fig. 1, we plot the profile of the metric function $f(r)$ and the horizon radius $r_{h}$ for various $\alpha$. From the left panel, one can see that there exists a threshold value of $\alpha=\alpha_{c}\simeq 2.56$ only above which $f(r)$ has zero points. This means that only when $\alpha\geq\alpha_{c}$, the metric represents a black hole and has two horizons when $\alpha>\alpha_{c}$. Moreover, as $r\rightarrow 0$ or $r\rightarrow\infty$, $f(r)\rightarrow 1$. From the right panel, one can see that the horizon radius $r_{h}$ increases with the increase of $\alpha$, and approaches to the Schwarzschild radius $r_{h}=2M$ in the limit of $\alpha\rightarrow\infty$.
@@ III Test charged scalar field
Now we consider a test charged scalar field $\Psi$ propagating in spacetime and search for the possible unstable modes under superradiance. Assume that the scalar field is minimally coupled to the electromagnetic field, so its dynamics is governed by the Klein-Gordon equation,
@@
To solve the radial equation (9) to obtain the frequency $\omega$ and the associated modes $u(r)$, we need to impose physical boundary conditions: the wave should be purely ingoing near the horizon, while bounded at infinity to trigger superradiant instability. With the tortoise coordinate defined as $dr_{\ast}=dr/f(r)$, the radial equation (9) can be cast into a Schordinger-like form
@@ V.1 Numerical strategy
where $\sigma=-i\tilde{\omega}/f^{\prime}(r_{h})$ and $c=M(\omega^{2}-k^{2})/k$. $N$ is the truncation order. The expansion series coefficients $u_{j}$ and $\tilde{u}_{j}$ are to be determined. The metric function $f(r)$ and electric potential $\Phi(r)$ can also be expanded in a similar form. Inserting these expansions into the radial equation (9), the higher-order coefficients $u_{j}$ ($\tilde{u}_{j}$) with $j\geq 1$ can be solved in terms of $u_{0}$ ($\tilde{u}_{0}$). Furthermore, considering that the radial equation is linear in $u(r)$, we can fix $u_{0}=1$. So at last there remain only one free coefficient $\tilde{u}_{0}$.
@@ VI Summary and discussions
The investigation presented here delves into the charged superradiant instability affecting a spherical regular black hole when subjected to scalar field perturbations. The spacetime model, as detailed in [48], is formulated using the gravitational-decoupling technique. This model smoothly transitions between Minkowski spacetime and the standard Schwarzschild black hole by introducing a parameter $\alpha$ that governs the deformation of the spacetime. Unlike other regular black holes such as the ABG black holes [15, 16], the deformation of this black hole from the standard Schwarzschild form shows an exponential decay along the radial direction. Through a combination of semi-analytical and numerical approaches, it has been observed that this particular black hole can exhibit charged superradiant instability within specific parameter regimes. Recent studies in Refs. [30, 31, 32] have already highlighted the presence of charged superradiant instability in ABG regular black holes. The findings of our current research further support the notion that regular black holes may generally manifest superradiant instability.

### v2

# Charged superradiant instability in a spherical regular black hole

@@ Abstract
We examine the stability of a spherically symmetric regular black hole when subjected to perturbations from a charged scalar field. This particular black hole is constructed by deforming the Minkowski spacetime. It has been observed that the charged superradiant instability arises only within a specific range of the deformation parameter, potentially resulting in an instability growth rate with a maximum magnitude of approximately $\text{Im}(M\omega)\sim 10^{-3}$. This growth rate significantly exceeds the instability identified in ABG black holes discussed in prior research, suggesting a notable timescale for detecting this phenomenon in astrophysical scenarios. Additionally, we conduct a thorough investigation into how the three parameters of the model influence the onset and intensity of the instability. Our analysis offers further insights into the possible emergence of this instability in spherically regular black holes and its association with the nonlinear effects of the electromagnetic field.
@@ I Introduction
Singularities are commonly considered non-physical within the classical theoretical framework due to their infinite energy density, prompting researchers to seek resolution through the inclusion of quantum effects in matter or gravity. While singularity theorems rely on strict conditions like the null energy condition for matter, allowing for the possibility of singularity-free black holes. To address the issue of non-physical singularities, a significant effort is being made by researchers to develop regular black hole solutions. This endeavor dates back to 1968 when Bardeen introduced a regular black hole model that adhered to the weak energy condition [14], circumventing the singularity theorems. Subsequently, Ayon-Beato and Garcia established a class of regular black holes known as ABG black holes [15, 16], featuring spherical charged configurations within the framework of Einstein-nonlinear electromagnetic fields. Further advancements led them to realize the Bardeen black hole as an exact solution within some Einstein-nonlinear-Maxwell theory [17]. These breakthroughs have inspired the development of various regular black hole models predominantly supported by nonlinear electromagnetic fields (see for example [18, 19, 20, 21, 22, 23] and references therein), which serve as an avenue to incorporate quantum effects in the electromagnetic field theory [24, 25, 26, 27]. Extensive investigations have been conducted on the dynamic behaviors, thermodynamic properties, and astrophysical implications of these black holes. The distinct observational effects exhibited by regular black holes compared to conventional ones have been a subject of interest, as indicated by recent studies and reviews on the topic [20, 21, 22, 28, 29, 30, 31].
@@
The authors of Refs. [32, 33, 34] recently conducted an extensive examination of the superradiance phenomenon in ABG black holes subjected to charged scalar field perturbations. Their findings suggest that within certain parameter ranges, ABG black holes display superradiant instability, a characteristic distinct from the well-established behavior of Reissner-Nordström (RN) black holes. The study of superradiance and its associated instability in black holes has been a focal point of research in black hole physics for a significant period. Superradiant instability typically manifests in rotating black holes [35, 36, 37, 38, 39]. However, for non-rotating, spherically symmetric RN black holes, superradiance can occur under charged scalar field perturbations but leading no instability [40, 41], unless the spacetime has a special boundary or an artificial boundary condition is imposed [42, 43, 44, 45, 46, 47]. The research outcomes in the Refs. [32, 33, 34] introduce a novel scenario: when accounting for the nonlinear effects of the electromagnetic field, spherically symmetric charged regular black holes may also exhibit superradiant instability. The superradiant instability of black holes holds significant astronomical observational implications. This mechanism results in the accumulation of scalar particles around the black hole, giving rise to observable effects such as alterations in gravitational waveforms and interference with the imaging of the black hole shadow [48, 49, 50, 51, 52, 53, 54]. These observable effects offer valuable insights and avenues for exploring exotic particles beyond the standard model and furthering our comprehension of the physical characteristics of black holes. See more discussions on this phenomenon and its astrophysical implications, please refer to the review [55] and references therein. Thus, the research findings in Refs. [32, 33, 34] not only unveil the superradiant instability of ABG black holes under specific circumstances but also present a potential observational approach for identifying and investigating the properties of regular black holes.
@@
The inquiry arises as to whether the phenomenon of superradiant instability is exclusive to ABG black holes or is a common occurrence among regular black holes. This study delves into this question by exploring a distinct class of regular black holes [56] generated through the “gravitational-decoupling” (GD) technique [57, 58], possessing unique characteristics. The spacetime of these black holes serves as an interpolation between Minkowski and Schwarzschild spacetime, characterized by a deformation parameter and potentially supported by a nonlinear electromagnetic field. Through a detailed analysis of the behavior of charged scalar field perturbations around such black holes, it is observed that the presence and strength of superradiant instability are closely linked to the deformation parameter. Remarkably, when the deformation parameter falls within a specific range, the intensity of superradiant instability can escalate significantly, surpassing that of ABG black holes. These findings provide evidence that regular black holes typically display superradiant instability.
@@
This paper is structured as follows: Section II provides a concise overview of the regular black hole under consideration. Section III delves into the equation of motion governing the tested charged scalar field. In Section IV, an examination of the parameter space is conducted to determine the permissible region for initiating the superradiant instability. Section V employs a numerical approach to identify the superradiant unstable modes. The concluding section encompasses a summary and discussions. The study employs natural units with $G=c=\hbar=4\pi\epsilon_{0}=1$ throughout, where $G,c,\hbar$ and $\epsilon_{0}$ are the Newtonian gravitational constant, the photon velocity in vacuum, the induced Planck constant and the vacuum permittivity, respectively.
@@ II The spacetime
where $M$ is the ADM mass of the black hole and $\alpha$ is a deformation parameter. It is constructed through the deformation of the Schwarzschild spacetime with the GD method [57, 58]. The metric represents a spacetime interpolating the Minkowski spacetime and the Schwarzschild black hole: In the limit of $\alpha\rightarrow 0$, it reduces to the Minkowski metric; while in the limit of $\alpha\rightarrow\infty$, it reduces to the standard Schwarzschild metric. It should also be noted that the deformation of this spacetime with respect to the Schwarzschild metric takes an exponential decaying form, so that the deformation mainly affects the small-$r$ region but leaves the large-$r$ region little affected. The spacetime has a de-Sitter core and is regular at $r=0$ which can be seen by calculating the typical curvature invariants, for example,
@@
It asymptotically takes a Yukawa-like form rather than a Maxwell-like form, which is very different from that of ABG black holes [15, 16]. So the “nonlinear electromagnetic field” may not be considered as the usual nonlinear extension of Maxwell’s theory simply, rather it is more likely associated with some kind of short-range interaction. The action from which to derive this solution is rather involved and not relevant to our present study, so we will not display it here. For more details, one can refer to Ref. [56].
The event horizon $r=r_{h}$ of the black hole is given by the largest root of $f(r)=0$, which depends on the value of $\alpha$. In Fig. 1, we plot the profile of the metric function $f(r)$ and the horizon radius $r_{h}$ for various $\alpha$. From the left panel, one can see that there exists a threshold value of $\alpha=\alpha_{c}\simeq 2.56$ only above which $f(r)$ has zero points. This means that only when $\alpha\geq\alpha_{c}$, the metric represents a black hole and has two horizons when $\alpha>\alpha_{c}$. Moreover, as $r\rightarrow 0$ or $r\rightarrow\infty$, $f(r)\rightarrow 1$. From the right panel, one can see that the horizon radius $r_{h}$ increases with the increase of $\alpha$, and approaches the Schwarzschild radius $r_{h}=2M$ in the limit of $\alpha\rightarrow\infty$.
@@ III Test charged scalar field
Now we consider a test charged scalar field $\Psi$ propagating in spacetime and search for the possible unstable modes under superradiance. Assume that the scalar field is minimally coupled to the electromagnetic field, so its dynamic is governed by the Klein-Gordon equation,
@@
To solve the radial equation (9) to obtain the frequency $\omega$ and the associated modes $u(r)$, we need to impose physical boundary conditions: the wave should be purely ingoing near the horizon, while bounded at infinity to trigger superradiant instability. With the tortoise coordinate defined as $dr_{\ast}=dr/f(r)$, the radial equation (9) can be cast into a Schrödinger-like form
@@ V.1 Numerical strategy
where $\sigma=-i\tilde{\omega}/f^{\prime}(r_{h})$ and $c=M(\omega^{2}-k^{2})/k$. $N$ is the truncation order. The expansion series coefficients $u_{j}$ and $\tilde{u}_{j}$ are to be determined. The metric function $f(r)$ and electric potential $\Phi(r)$ can also be expanded in a similar form. Inserting these expansions into the radial equation (9), the higher-order coefficients $u_{j}$ ($\tilde{u}_{j}$) with $j\geq 1$ can be solved in terms of $u_{0}$ ($\tilde{u}_{0}$). Furthermore, considering that the radial equation is linear in $u(r)$, we can fix $u_{0}=1$. So at last there remains only one free coefficient $\tilde{u}_{0}$.
@@ VI Summary and discussions
The investigation presented here delves into the charged superradiant instability affecting a spherical regular black hole when subjected to scalar field perturbations. The spacetime model, as detailed in [56], is formulated using the GD technique. This model smoothly transitions between Minkowski spacetime and the standard Schwarzschild black hole by introducing a parameter $\alpha$ that governs the deformation of the spacetime. Unlike other regular black holes such as the ABG black holes [15, 16], the deformation of this black hole from the standard Schwarzschild form shows an exponential decay along the radial direction. Through a combination of semi-analytical and numerical approaches, it has been observed that this particular black hole can exhibit charged superradiant instability within specific parameter regimes. Recent studies in Refs. [32, 33, 34] have already highlighted the presence of charged superradiant instability in ABG regular black holes. The findings of our current research further support the notion that regular black holes may generally manifest superradiant instability.